OpenStudy (anonymous):

Find a function y of x such that 9yy′=x and y(9)=10.

OpenStudy (anonymous):

\[9y~\frac{ dy }{ dx }=x\] separate the variables and integrate.

OpenStudy (anonymous):

9 y dy=x dx\[\int\limits 9y dy=\int\limits x dx\] \[9\frac{ y^2 }{ 2 }=\frac{ x^2 }{ 2}+c\] when x=9,y=10 find c write the solution.